# On Representation Complexity of Model-based and Model-free Reinforcement Learning

### v1

@@ B.2 Proof of Theorem 1
Lemma 1 (Property of value function in unconditioned Majority MDPs).
@@
For the control bits to traverse from $\mathbf{0}_{b}$ to $\mathbf{1}_{b}$, it takes $12^{b}-$ 1 time steps. Indeed, by Definition 7, $f^{(i)}(\mathbf{0}_{b})\neq\mathbf{1}_{b}$ for any $i<2^{b}-1$ and $f^{(2^{b}-1)}(\mathbf{0}_{b})=\mathbf{1}_{b}$ (otherwise, |$\left|\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}\right|<2^{b}-1$ and as a result $\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}$ can not traverse $\{0,1\}^{b}\backslash\{\mathbf{0}_{b}\}$). Thus, this corresponds to at least $12^{b}-$ 1 time steps in which action $a=0$ is played.
@@
$\pi^{*}(s)=\begin{cases}1,&~{}\text{if}~{}s_{b+i}\neq s_{\mathrm{reward},i},~{}\text{where}~{}i=\sum_{j=1}^{b}2^{b-j}\cdot s_{j}\\ 0,&~{}\text{otherwise}.\end{cases}$
@@
By Claim 2.13 in Arora and Barak [2009], any Boolean function can be computed by a CNF formula, i.e., Boolean circuits of the form
@@
Now we return to the proof of Theorem 1.
@@
where $\oplus$ is the XOR gate, $g:\{0,1\}^{b+1}\to\{0,1\}^{b}$ such that $g(x_{1},\dots,x_{b},a)_{i}=(f(x_{1},\dots,x_{b})_{i}\wedge\neg a)\vee(x_{i}\wedge a)$ for all $i\in[b]$, and $\delta_{k}(x_{1},\dots,x_{b})=\mathbbm{1}(k=\sum_{j=1}^{b}2^{b-j}\cdot x_{j})$. We can verify that $C_{\mathrm{model}}=\mathbb{T}$. By Claim B.2, Claim B.3, and Lemma 2, the XOR gate, the control function $f$, and the gate $\delta_{k}$ can all be implemented by binary circuits with polynomial size and constant depth. As a result of Claim B.1, $C_{\mathrm{model}}$ also has polynomial size and constant depth.
@@
Finally, we show that the value function can not be computed by a circuit with constant depth and polynomial size. By Lemma 1, we have
@@
outputs the majority function in the first bit. By Proposition A.2, $C_{\mathsf{MAJORITY}}$ also has polynomial size and constant depth, which contradicts Proposition 2. This means that $V^{*}$ can not be computed by circuits with polynomial size and constant depth. Notice that $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)=\max\{Q^{*}(s,0),Q^{*}(s,1)\}$. By Proposition A.3 and Claim B.1, we conclude that the optimal $Q$-function can not be computed by circuits with polynomial size and constant depth. ∎
B.3 Proof of Theorem 2
@@
$\mathbb{P}\left(C(x_{1},\dots,x_{n})=1\right)=$ $~{}1-\prod_{i=1}^{n}\mathbb{P}\left(\left(\bigwedge_{j=1}^{k_{i}}X_{i,j}\right)=0\right)$
$=$ $~{}1-\prod_{i=1}^{n}(1-2^{-k_{i}})$
$\geq$ $~{}1-e^{-\Omega(m)}.$
Lemma 4 (Property of value function in conditioned Majority MDPs).
In an $n$-bits majority MDP with reward state $s_{\mathrm{reward}}$, control function $f$ and condition $C$, if there exists a set $A\subset[n]$ such that $|A|=\lceil n/2\rceil=2^{b-1}-1$ and $C(x)=1$ holds for any binary string $x\in\{0,1\}^{n}$ satisfying $x_{i}=s_{\mathrm{reward},i},\forall i\notin A$, then the value function $V^{*}$ over $\{s\in\{0,1\}^{b+n}:s[c]=\mathbf{0}_{b},s_{b+i}=s_{\mathrm{reward},i}~{}(\forall i\notin A)\}$ is given by the following:
@@
For the control bits to travel from $\mathbf{0}_{b}$ to $\mathbf{1}_{b}$, it takes $12^{b}-$ 1 time steps. Indeed, by Definition 7, $f^{(i)}(\mathbf{0}_{b})\neq\mathbf{1}_{b}$ for any $i<2^{b}-1$ and $f^{(2^{b}-1)}(\mathbf{0}_{b})\neq\mathbf{1}_{b}$ (otherwise, |$\left|\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}\right|<2^{b}-1$ and as a result $\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}$ can not traverse $\{0,1\}^{b}\backslash\{\mathbf{0}_{b}\}$). Thus, this corresponds to at least $12^{b}-$ 1 time steps in which action $a=0$ is played.
@@
$\pi^{*}(s)=\begin{cases}1,&~{}\text{if}~{}s_{b+i}\neq s_{\mathrm{reward},i},~{}\text{where}~{}i=\sum_{j=1}^{b}2^{b-j}\cdot s_{j}\\ 0,&~{}\text{otherwise}.\end{cases}$
@@
Now we return to the proof of Theorem 2.
@@
where $\oplus$ is the XOR gate, $g:\{0,1\}^{b+1}\to\{0,1\}^{b}$ such that $g(x_{1},\dots,x_{b},a)_{i}=(f(x_{1},\dots,x_{b})_{i}\wedge\neg a)\vee(x_{i}\wedge a)$ for all $i\in[b]$, and $\delta_{k}(x_{1},\dots,x_{b})=\mathbbm{1}(k=\sum_{j=1}^{b}2^{b-j}\cdot x_{j})$. We can verify that $C_{\mathrm{model}}=\mathbb{T}$. By Claim B.2, Claim B.3, and Lemma 2, the XOR gate, the control function $f$, the function $C(s[r])$, and the gate $\delta_{k}$ can all be implemented by binary circuits with polynomial size and constant depth. As a result of Claim B.1, $C_{\mathrm{model}}$ also has polynomial size and constant depth.
@@
Finally, we show that with high probability, the value function can not be computed by a circuit with constant depth and polynomial size. By Lemma 3, with probability at least $1-e^{-\Omega(m)}$, there exists a set $A\subset[n]$ such that $|A|\geq n/2$ and $C(x)=1$ holds for any binary string $x\in\{0,1\}^{n}$ satisfying $x_{i}=s_{\mathrm{reward},i},\forall i\notin A$. We can reduce the size of $A$ by deleting some elements to make $|A|=\lceil n/2\rceil=2^{b-1}-1$ and the above property still holds. Denote $A=\{a(1)<\dots<a(L)\}$ where $a:[L]\to[n]$. Define $h:\{0,1\}^{L}\to\{0,1\}^{b}$
$h\left((x_{1},\dots,x_{L})\right)=n+1-V^{*}\left((\mathbf{0}_{b},s^{\prime})\right),~{}\text{where}~{}s^{\prime}_{j}=\begin{cases}s_{\mathrm{reward},j},&~{}j\notin A\\ \neg x_{a^{-1}(j)}\oplus s_{\mathrm{reward},j},&~{}j\in A\end{cases}.$
Due to Lemma 4, $h\left(x\right)$ can be represented in binary form with $(b-1)$ bits, and the first (most significant) bit of $h(x)$ is the majority function of $x$. If there exists a circuit $C_{\mathrm{value}}=V^{*}$ with polynomial size and constant depth, then the circuit defined by
@@
where $a^{-1}(i)=i$ if $i\notin A$, outputs $h$. By Proposition A.2, $C_{\mathsf{MAJORITY}}$ also has polynomial size and constant depth, which contradicts Proposition A.1. This means that $V^{*}$ can not be computed by circuits with polynomial size and constant depth. Notice that $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)=\max\{Q^{*}(s,0),Q^{*}(s,1)\}$. By Proposition A.3 and Claim B.1, we conclude that the optimal $Q$-function can not be computed by circuits with polynomial size and constant depth. ∎
@@ Appendix C Experiment Details
Table 1 shows the parameters used in SAC training to learn the optimal $Q$-function. Table 2 shows parameters for fitting neural networks to the value, reward, and transition functions.
@@
Optimizer Adam [Kingma and Ba, 2014]
@@
Optimizer Adam [Kingma and Ba, 2014]

### v2

@@ B.1 Useful results
Lemma B.4 (Property of value function in unconditioned Majority MDPs).
@@
For the control bits to traverse from $\mathbf{0}_{b}$ to $\mathbf{1}_{b}$, it takes $12^{b}-$ 1 time steps. Indeed, by Definition 3.2, $f^{(i)}(\mathbf{0}_{b})\neq\mathbf{1}_{b}$ for any $i<2^{b}-1$ and $f^{(2^{b}-1)}(\mathbf{0}_{b})=\mathbf{1}_{b}$ (otherwise, |$\left|\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}\right|<2^{b}-1$ and as a result $\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}$ can not traverse $\{0,1\}^{b}\backslash\{\mathbf{0}_{b}\}$). Thus, this corresponds to at least $12^{b}-$ 1 time steps in which action $a=0$ is played.
@@
$\pi^{*}(s)=\begin{cases}1,&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ s_{b+i}\neq s_{\mathrm{reward},i},\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ i=\sum_{j=1}^{b}2^{b-j}\cdot s_{j}\\ 0,&\leavevmode\nobreak\ \text{otherwise}.\end{cases}$
@@
By Claim 2.13 in Arora & Barak (2009), any Boolean function can be computed by a CNF formula, i.e., Boolean circuits of the form
@@
Now we return to the proof of Theorem 3.7.
@@
where $\oplus$ is the XOR gate, $g:\{0,1\}^{b+1}\to\{0,1\}^{b}$ such that $g(x_{1},\dots,x_{b},a)_{i}=(f(x_{1},\dots,x_{b})_{i}\wedge\neg a)\vee(x_{i}\wedge a)$ for all $i\in[b]$, and $\delta_{k}(x_{1},\dots,x_{b})=\mathbbm{1}(k=\sum_{j=1}^{b}2^{b-j}\cdot x_{j})$. We can verify that $C_{\mathrm{model}}=\mathbb{T}$. By Claim B.2, Claim B.3, and Lemma B.5, the XOR gate, the control function $f$, and the gate $\delta_{k}$ can all be implemented by binary circuits with polynomial size and constant depth. As a result of Claim B.1, $C_{\mathrm{model}}$ also has polynomial size and constant depth.
@@
Finally, we show that the value function can not be computed by a circuit with constant depth and polynomial size. By Lemma B.4, we have
@@
outputs the majority function in the first bit. By Proposition A.4, $C_{\mathsf{MAJORITY}}$ also has polynomial size and constant depth, which contradicts Proposition 3.8. This means that $V^{*}$ can not be computed by circuits with polynomial size and constant depth. Notice that $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)=\max\{Q^{*}(s,0),Q^{*}(s,1)\}$. By Proposition A.6 and B.1, we conclude that the optimal $Q$-function can not be computed by circuits with polynomial size and constant depth. ∎
B.3 Proof of Theorem 3.8
@@
$\mathbb{P}\left(C(x_{1},\dots,x_{n})=1\right)=$ $\leavevmode\nobreak\ 1-\prod_{i=1}^{n}\mathbb{P}\left(\left(\bigwedge_{j=1}^{k_{i}}X_{i,j}\right)=0\right)$
$=$ $\leavevmode\nobreak\ 1-\prod_{i=1}^{n}(1-2^{-k_{i}})$
$\geq$ $\leavevmode\nobreak\ 1-e^{-\Omega(m)}.$
Lemma B.7 (Property of value function in conditioned Majority MDPs).
In an $n$-bits majority MDP with reward state $s_{\mathrm{reward}}$, control function $f$ and condition $C$, if there exists a set $A\subset[n]$ such that $|A|=\lceil n/2\rceil=2^{b-1}-1$ and $C(x)=1$ holds for any binary string $x\in\{0,1\}^{n}$ satisfying $x_{i}=s_{\mathrm{reward},i},\forall i\notin A$, then the value function $V^{*}$ over $\{s\in\{0,1\}^{b+n}:s[c]=\mathbf{0}_{b},s_{b+i}=s_{\mathrm{reward},i}\leavevmode\nobreak\ (\forall i\notin A)\}$ is given by the following:
@@
For the control bits to travel from $\mathbf{0}_{b}$ to $\mathbf{1}_{b}$, it takes $12^{b}-$ 1 time steps. Indeed, by Definition 3.2, $f^{(i)}(\mathbf{0}_{b})\neq\mathbf{1}_{b}$ for any $i<2^{b}-1$ and $f^{(2^{b}-1)}(\mathbf{0}_{b})\neq\mathbf{1}_{b}$ (otherwise, |$\left|\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}\right|<2^{b}-1$ and as a result $\left\{f^{(k)}(\mathbf{0}_{b}):k=1,2,\dots,2^{b}-1\right\}$ can not traverse $\{0,1\}^{b}\backslash\{\mathbf{0}_{b}\}$). Thus, this corresponds to at least $12^{b}-$ 1 time steps in which action $a=0$ is played.
@@
$\pi^{*}(s)=\begin{cases}1,&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ s_{b+i}\neq s_{\mathrm{reward},i},\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ i=\sum_{j=1}^{b}2^{b-j}\cdot s_{j}\\ 0,&\leavevmode\nobreak\ \text{otherwise}.\end{cases}$
@@
Now we return to the proof of Theorem 3.8.
@@
where $\oplus$ is the XOR gate, $g:\{0,1\}^{b+1}\to\{0,1\}^{b}$ such that $g(x_{1},\dots,x_{b},a)_{i}=(f(x_{1},\dots,x_{b})_{i}\wedge\neg a)\vee(x_{i}\wedge a)$ for all $i\in[b]$, and $\delta_{k}(x_{1},\dots,x_{b})=\mathbbm{1}(k=\sum_{j=1}^{b}2^{b-j}\cdot x_{j})$. We can verify that $C_{\mathrm{model}}=\mathbb{T}$. By Claim B.2, Claim B.3, and Lemma B.5, the XOR gate, the control function $f$, the function $C(s[r])$, and the gate $\delta_{k}$ can all be implemented by binary circuits with polynomial size and constant depth. As a result of Claim B.1, $C_{\mathrm{model}}$ also has polynomial size and constant depth.
@@
Finally, we show that with high probability, the value function can not be computed by a circuit with constant depth and polynomial size. By Lemma B.6, with probability at least $1-e^{-\Omega(m)}$, there exists a set $A\subset[n]$ such that $|A|\geq n/2$ and $C(x)=1$ holds for any binary string $x\in\{0,1\}^{n}$ satisfying $x_{i}=s_{\mathrm{reward},i},\forall i\notin A$. We can reduce the size of $A$ by deleting some elements to make $|A|=\lceil n/2\rceil=2^{b-1}-1$ and the above property still holds. Denote $A=\{a(1)<\dots<a(L)\}$ where $a:[L]\to[n]$. Define $h:\{0,1\}^{L}\to\{0,1\}^{b}$
$h\left((x_{1},\dots,x_{L})\right)=n+1-V^{*}\left((\mathbf{0}_{b},s^{\prime})\right),\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ s^{\prime}_{j}=\begin{cases}s_{\mathrm{reward},j},&\leavevmode\nobreak\ j\notin A\\ \neg x_{a^{-1}(j)}\oplus s_{\mathrm{reward},j},&\leavevmode\nobreak\ j\in A\end{cases}.$
Due to Lemma B.7, $h\left(x\right)$ can be represented in binary form with $(b-1)$ bits, and the first (most significant) bit of $h(x)$ is the majority function of $x$. If there exists a circuit $C_{\mathrm{value}}=V^{*}$ with polynomial size and constant depth, then the circuit defined by
@@
where $a^{-1}(i)=i$ if $i\notin A$, outputs $h$. By Proposition A.4, $C_{\mathsf{MAJORITY}}$ also has polynomial size and constant depth, which contradicts Proposition A.2. This means that $V^{*}$ can not be computed by circuits with polynomial size and constant depth. Notice that $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)=\max\{Q^{*}(s,0),Q^{*}(s,1)\}$. By Proposition A.6 and B.1, we conclude that the optimal $Q$-function can not be computed by circuits with polynomial size and constant depth. ∎
@@ Appendix C Experiment Details
Table 1 shows the parameters used in SAC training to learn the optimal $Q$-function in Section 4. Table 2 shows parameters for fitting neural networks to the value, reward, and transition functions in Section 4.
@@
Optimizer Adam (Kingma & Ba, 2014)
@@
Optimizer Adam (Kingma & Ba, 2014)
@@
Appendix D Additional Experiment Results
Under the same experiment settings in Section 4, we conduct further experiments using varying neural network depth and width to approximate the the optimal $Q$-functions, reward functions, and transition functions. The results are shown in Figure 4-Figure 7. Furthermore, since the optimal $Q$-function calculated by the critic in the SAC algorithm might be biased due to under-estimation since clipped double Q learning is adopted, we also use critics of the Actor-Critic algorithm (Section D.1) and Monte Carlo simulation to evaluate the $Q$-function of the learned actor (Section D.2) to obtain an unbiased estimation of the optimal $Q$-function.
Figure 4: Approximation errors of the optimal $Q$-functions, reward functions, and transition functions in MuJoCo environments, using $2$-(hidden) layer neural networks with width $128$. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors.
Figure 5: Approximation errors of the optimal $Q$-functions, reward functions, and transition functions in MuJoCo environments, using $1$-(hidden) layer neural networks with width $16$. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors.
Figure 6: Approximation errors of the optimal $Q$-functions, reward functions, and transition functions in MuJoCo environments, using $3$-(hidden) layer neural networks with width $64$. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors.
Figure 7: Approximation errors of the optimal $Q$-functions, reward functions, and transition functions in MuJoCo environments, using $2$-(hidden) layer neural networks with width $64$. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors.
D.1 Fitting Q-values learned by Actor-Critic
In addition to approximating the Q-functions learned from SAC, we also fit the value functions learned in Actor-Critic algorithm. The result, displayed in Figure 8, exhibits that the value functions are more difficult to approximate than the reward and the transition functions.
Figure 8: Approximation errors of the optimal value functions (learned by actor-critic), reward functions, and transition functions in MuJoCo environments, using $3$-(hidden) layer neural networks with width $64$. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors.
D.2 Fitting Q-values from Monte-Carlo sampling
We also consider Monte-Carlo estimate of the Q-values of the actors learned by SAC. In Figure 9, we visualize the approximation errors (in log scale), which confirms the phenomenon found in Section 4.
Figure 9: Approximation errors of the optimal $Q$-functions, reward functions, and transition functions in MuJoCo environments, using $2$-(hidden) layer neural networks with width $128$. For each state, we take $3$ Monte-Carlo sampling and use the average discounted sum of rewards as the Q-value estimate. The approximation errors are shown in the log scale. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors.
Appendix E Extensions to Stochastic MDP and POMDP
In this section, we briefly discuss possible approaches of generalizing our frameworks to stochastic MDP and partially observable MDP (POMDP).
Stochastic MDP.
Although we study the deterministic transition in this paper, it can also be extended to stochastic MDPs. In that case, $P(s,a,s\textquoteright)$ ’ ) represents the probability of reaching state $s\textquoteright$ ’ when choosing action $a$ on state $s$. Since the probability is represented with bounded precision in machine (e.g., 64-bit floating point), the output of $P(s,a,s\textquoteright)$ ’ ) can also be represented by a $\{0,1\}$-string.
Now we discuss a natural extension from the deterministic transition kernel to the stochastic transition kernel of our majority MDP. After applying $a=1$, if $C(s[r])=1$ (i.e., the condition is satisfied), instead of deterministically flipping the $s[c]$-th bit of representation bits, in the stochastic version, the $s[c]$-th bit will be flipped with probability $1/2$, and nothing happens otherwise. This resembles the real-world scenario where an action might fail with some failure probability. The representation complexity of the transition kernel $P$ remains polynomial. For the optimal $Q$ function, if we change $H=2^{b}+2n$, since in expectation, each (effective) flipping will cost two steps, the representation complexity of the $Q$-function is still lower bounded by the majority function, which is exponentially large. Note that our deterministic majority MDP can be viewed as an embedded chain of a more ‘lazy’ MDP, and the failure probability can also be arbitrary. Hence, our result (Theorems 3.8 and 3.7) can be generalized to this natural class of stochastic MDPs.
POMDP.
For POMDP, we can additionally assume a function $O(s)=o$ (or $O(s,a)=o$), which maps a state $s$ (or a state-action pair $(s,a)$) to a corresponding observation $o$. One possible choice of $O(s)$ could be a substring of the state $s$, which still has low representation complexity. In POMDPs, the $Q$-function is often defined on a sequence of past observations or certain belief states that reflect the posterior distribution of the hidden state. As a result, the $Q$-functions in POMDPs often have higher complexity. Since the $Q$-function is proven to possess high circuit complexity for the majority MDP, the representation complexity for value-based quantities of a partially observable version could be even higher.